OpenStudy (anonymous):

Simplify 3(-3/2)^-3 -9^-1 + 9^0

OpenStudy (anonymous):

3(-3/2)^-3 -9^-1 + 9^0 =3(-2/3)^3 -1/9 +1 =3(-8/27) -1/9 +1 =-8/9 -1/9 +1 =-9/9 +1 =1-1=0 ans

OpenStudy (anonymous):

Thank you!